# Semantics-aware Motion Retargeting with Vision-Language Models

### v1

@@ 3.3 Semantics & Geometry Fine-tuning
Prompt design. Since the vision-language model has the capability to extract rich information from images, it is possible that the extracted features might contain redundant details, such as the appearance of the character. To guide the vision-language model to obtain semantic embedding relevant to character motions, we adopt a guiding visual question answering approach for motion semantics extraction, as depicted in Fig. 3. We believe that there is a strong correlation between motion semantics and hand movements. To acquire a more comprehensive description of the motion, we initially provide a guiding question to BLIP-2: “Where are the hands of the character?”. Subsequently, we introduce a new question and combine it with the first answer as the input to BLIP-2: “[The answers to the first question generated by the vision-language model] What is the character in the image doing?”. For more details, please refer to the supplementary materials.
@@ 4.1 Settings
Datasets. We train and evaluate our method on the Mixamo dataset [1], an extensive repository of animations performed by various 3D virtual characters with distinct skeletons and geometry shapes. The training set we use to pre-train our skeleton aware module is the same as that used in [2], which contains 1646 motions performed by 7 characters. It’s important to note that the Mixamo dataset does not provide clean ground truth data, since many of the motion sequences suffer from interpenetration issues and semantic information loss. To mitigate this, we have carefully selected a subset of motion sequences that are both semantically clean and free of interpenetration issues for fine-tuning and testing. Our fine-tuning process involves retargeting 15 clean motions including 3127 frames, originally performed by 3 source characters, namely “Y Bot”, “X Bot”, and “Ortiz”, onto 3 target characters, including “Aj”, “Kaya”, and “Mousey”. Then we evaluate the performance of our model on the task of retargeting 30 additional motions that are previously unseen in the training set and fine-tuning sets. More details could be found in the supplementary materials.
@@ 4.2 Comparison with State of the Arts
| Method | MSE | MSE | Pen.% | ITM | FID | SCL |
| SMT | 0.248 | 0.129 | 8.37 | 0.586 | 7.727 | 0.769 |
| SMT | 7.798 | 7.083 | 0.44 | 0.432 | 56.53 | 13.29 |
| SMT | 0.335 | 0.288 | 5.36 | 0.658 | 2.826 | 0.266 |
| SMT | 0.439 | 0.368 | 1.22 | 0.597 | 7.241 | 0.583 |
| SMT | 5.418 | 4.576 | 4.41 | 0.552 | 78.46 | 18.96 |
| SMT | 0.739 | 0.517 | 4.56 | 0.668 | 2.497 | 0.191 |
| SMT | 0.284 | 0.229 | 3.50 | 0.680 | 0.436 | 0.143 |
@@ 4.3 Ablation Studies
Skeleton-aware pre-training. The proposed method can be divided into two stage: pre-training and fine-tuning. To illustrate the importance of skeleton-aware pre-training, we evaluate the network trained with only the semantics consistency loss and the interpenetration loss in Tab. 2, denoted as SMT. The network trained without skeleton-aware pre-training performs worst in MSE and semantics preservation. A reasonable explanation is that the semantics consistency loss is highly non-linear, so it is important to pre-train the network at the skeletal level to provide better initial values. We also visualize qualitative results in Fig. 5.
Semantics & geometry fine-tuning. We also conduct ablation study to illustrate the importance of semantics and geometry fine-tuning in Tab. 2. We first evaluate the performance of the skeleton-aware model without fine-tuning, denoted as SMT. Though it reaches the best global position MSE, it suffers from interpenetration and semantic information loss because of the low-quality motion data provided by Mixamo. We next evaluate the network fine-tuned with only the interpenetration loss, denoted as SMT. This version results in a significant boost in terms of penetration rate. However, the gradient of interpenetration loss is only relevant with the face normals of the geometry mesh without considering the semantic information conveyed in the motion. It indicates the importance of the semantic consistency loss that makes the network reach a better balance between interpenetration and semantics. We also try to train the network with all loss functions in one stage, denoted as SMT. However, it is challenging for the model to acquire general knowledge of interpenetration and semantics that is suitable for every character with limited data. Therefore, training the model with skeleton-aware pre-training and fine-tuning it with semantics consistency and geometry constraints for each target character remains a more reasonable and data-efficient strategy.
Latent semantic embedding. The vision-language model used for semantic extraction can be divided into three parts: the image encoder from CLIP [22], the querying transformer and the large language model. In Tab. 2, we compare the feature outputted by the image encoder, the querying transformer and the encoder of the large language model, denoted as SMT, SMT, and SMT, respectively. The results show that the image feature performs worse since it is greatly affected by the appearance of the character. It indicates that with the help of the large language model, the semantic representation better focuses on the semantic meaning of the motion instead of the character’s visual appearance. Therefore, the encoder output of the large language model is more suitable for semantic embedding. More details can be found in the supplementary materials.
@@ C Implementation Details
| Method | MSE | MSE | Pen.% | ITM | FMD | SCL |
| SMT | 5.418 | 4.576 | 4.41 | 0.552 | 78.46 | 18.96 |
| SMT | 0.739 | 0.517 | 4.56 | 0.658 | 2.497 | 0.191 |
| SMT | 0.284 | 0.229 | 3.50 | 0.680 | 0.436 | 0.143 |
@@ E Ablation Study
Latent semantic embedding. We validate features from different levels: the output of the image encoder, the output of the querying transformer, the output of the large language model encoder. We visualize the features of three motions including Waving, Pointing and Salute with three source characters, including Y Bot, X Bot, Ortiz, using T-SNE [25] dimensionality reduction technique. The Fig. 11 shows that the features of the image encoder are clustered around characters rather than motions which indicates that the features contains more information on the appearance of the characters, while the features of the large language model encoder are clustered around motions and contains mainly motion semantics. We further use these features as semantic embedding to fine-tune our model and compute evaluation metrics in Tab. 4. The metrics and qualitative comparison indicates that appearance of character may lead to meaningless gradient for the model resulting in unnatural retargeted motion. Moreover, compared with the image features from CLIP, the features of the querying transformer and the large language model encoder focus more on relevant semantics with the help of guiding questions.

### v3

@@ 3.3 Semantics & Geometry Fine-tuning
Prompt design. Since the vision-language model has the capability to extract rich information from images, it is possible that the extracted features might contain redundant details, such as the appearance of the character. To guide the vision-language model to obtain semantic embedding relevant to character motions, we adopt a guiding visual question answering approach for motion semantics extraction, as depicted in Fig. 3. We believe that there is a strong correlation between motion semantics and hand movements. To acquire a more comprehensive description of the motion, we initially provide a guiding question to BLIP-2: “Where are the hands of the character?". Subsequently, we introduce a new question and combine it with the first answer as the input to BLIP-2: “[The answers to the first question generated by the vision-language model] What is the character in the image doing?". For more details, please refer to the supplementary materials.
@@ 4.1 Settings
Datasets. We train and evaluate our method on the Mixamo dataset [1], an extensive repository of animations performed by various 3D virtual characters with distinct skeletons and geometry shapes. The training set we use to pre-train our skeleton aware module is the same as that used in [2], which contains 1646 motions performed by 7 characters. It’s important to note that the Mixamo dataset does not provide clean ground truth data, since many of the motion sequences suffer from interpenetration issues and semantic information loss. To mitigate this, we have carefully selected a subset of motion sequences that are both semantically clean and free of interpenetration issues for fine-tuning and testing. Our fine-tuning process involves retargeting 15 clean motions including 3127 frames, originally performed by 3 source characters, namely “Y Bot", “X Bot", and “Ortiz", onto 3 target characters, including “Aj", “Kaya", and “Mousey". Then we evaluate the performance of our model on the task of retargeting 30 additional motions that are previously unseen in the training set and fine-tuning sets. More details could be found in the supplementary materials.
@@ 4.2 Comparison with State of the Arts
| Method | MSE | MSElc | Pen.% | ITM | FID | SCL |
| SMTtws | 0.248 | 0.129 | 8.37 | 0.586 | 7.727 | 0.769 |
| SMTtwf | 7.798 | 7.083 | 0.44 | 0.432 | 56.53 | 13.29 |
| SMTtwa | 0.335 | 0.288 | 5.36 | 0.658 | 2.826 | 0.266 |
| SMTfwp | 0.439 | 0.368 | 1.22 | 0.597 | 7.241 | 0.583 |
| SMTfwi | 5.418 | 4.576 | 4.41 | 0.552 | 78.46 | 18.96 |
| SMTfwq | 0.739 | 0.517 | 4.56 | 0.668 | 2.497 | 0.191 |
| SMTOurs | 0.284 | 0.229 | 3.50 | 0.680 | 0.436 | 0.143 |
@@ 4.3 Ablation Studies
Skeleton-aware pre-training. The proposed method can be divided into two stage: pre-training and fine-tuning. To illustrate the importance of skeleton-aware pre-training, we evaluate the network trained with only the semantics consistency loss and the interpenetration loss in Tab. 2, denoted as SMTtwf. The network trained without skeleton-aware pre-training performs worst in MSE and semantics preservation. A reasonable explanation is that the semantics consistency loss is highly non-linear, so it is important to pre-train the network at the skeletal level to provide better initial values. We also visualize qualitative results in Fig. 5.
Semantics & geometry fine-tuning. We also conduct ablation study to illustrate the importance of semantics and geometry fine-tuning in Tab. 2. We first evaluate the performance of the skeleton-aware model without fine-tuning, denoted as SMTtws. Though it reaches the best global position MSE, it suffers from interpenetration and semantic information loss because of the low-quality motion data provided by Mixamo. We next evaluate the network fine-tuned with only the interpenetration loss, denoted as SMTfwp. This version results in a significant boost in terms of penetration rate. However, the gradient of interpenetration loss is only relevant with the face normals of the geometry mesh without considering the semantic information conveyed in the motion. It indicates the importance of the semantic consistency loss that makes the network reach a better balance between interpenetration and semantics. We also try to train the network with all loss functions in one stage, denoted as SMTtwa. However, it is challenging for the model to acquire general knowledge of interpenetration and semantics that is suitable for every character with limited data. Therefore, training the model with skeleton-aware pre-training and fine-tuning it with semantics consistency and geometry constraints for each target character remains a more reasonable and data-efficient strategy.
Latent semantic embedding. The vision-language model used for semantic extraction can be divided into three parts: the image encoder from CLIP [22], the querying transformer and the large language model. In Tab. 2, we compare the feature outputted by the image encoder, the querying transformer and the encoder of the large language model, denoted as SMTfwi, SMTfwq, and SMTOurs, respectively. The results show that the image feature performs worse since it is greatly affected by the appearance of the character. It indicates that with the help of the large language model, the semantic representation better focuses on the semantic meaning of the motion instead of the character’s visual appearance. Therefore, the encoder output of the large language model is more suitable for semantic embedding. More details can be found in the supplementary materials.
@@ 5 Conclusions
Acknowledgements. This work was supported by the National Nature Science Foundation of China under Grant 62173293.
@@ C Implementation Details
| Method | MSE | MSElc | Pen.% | ITM | FMD | SCL |
| SMTfwi | 5.418 | 4.576 | 4.41 | 0.552 | 78.46 | 18.96 |
| SMTfwq | 0.739 | 0.517 | 4.56 | 0.658 | 2.497 | 0.191 |
| SMTOurs | 0.284 | 0.229 | 3.50 | 0.680 | 0.436 | 0.143 |
@@ E Ablation Study
Latent semantic embedding. We validate features from different levels: the output of the image encoder, the output of the querying transformer, the output of the large language model encoder. We visualize the features of three motions including Waving, Pointing and Salute with three source characters, including Y Bot, X Bot, Ortiz, using T-SNE [25] dimensionality reduction technique. The Fig. 11 shows that the features of the image encoder are clustered around characters rather than motions which indicates that the features contains more information on the appearance of the characters, while the features of the large language model encoder are clustered around motions and contains mainly motion semantics. We further use these features as semantic embedding to fine-tune our model and compute evaluation metrics in Tab. 4. The metrics and qualitative comparison indicates that appearance of character may lead to meaningless gradient for the model resulting in unnatural retargeted motion. Moreover, compared with the image features from CLIP, the features of the querying transformer and the large language model encoder focus more on relevant semantics with the help of guiding questions.Further more, We visualize the semantic embeddings outputed by VLM in Figure 12. The visualization indicates that VLM has captured rich motion semantics information, regardless of the appearance of the character, and is capable of guiding motion semantics preservation.
@@
Performance of different numbers of views. We have conducted a new ablation in Table 6. The performance decreases to some extent with a single view due to depth information loss. Three perspectives, including the front, left and right view, can achieve fairly good results. Adding additional viewpoints after the third improves the outcome, but at a slower rate.
Ablation study of skeleton network. We have conducted a new ablation experiment in Table 6. The results show that the semantic module is applicable for different backbones and improves motion semantics preservation. The MSE metric has increased because the ground truth data in Mixamo dataset are not clean and suffer from interpenetration issues and semantic information loss [27].
| Number of Views | MSE | MSElc | Pen.% | ITM | FID | SCL |
| 1(front) | 0.262 | 0.186 | 3.51 | 0.630 | 5.715 | 0.499 |
| 2(left right diagonal) | 0.274 | 0.195 | 3.49 | 0.651 | 2.849 | 0.277 |
| 3(front,left,right) | 0.284 | 0.229 | 3.50 | 0.680 | 0.436 | 0.143 |
| 5(2+3) | 0.286 | 0.233 | 3.50 | 0.681 | 0.433 | 0.141 |
| Method | MSE | MSElc | Pen.% | ITM | FID | SCL |
| NKN [26] | 0.326 | 0.231 | 8.71 | 0.575 | 27.79 | 1.414 |
| NKN [26] + VLM | 0.392 | 0.308 | 4.44 | 0.665 | 2.687 | 0.223 |
| SAN [2] | 0.435 | 0.255 | 9.74 | 0.561 | 28.33 | 1.448 |
| SAN [2] + VLM | 0.481 | 0.339 | 5.08 | 0.659 | 2.798 | 0.258 |